OpenStudy (anonymous):

Gold in the form of solid cylinders will be molded into solid blocks. Each cylinder has a diameter of 6.19 cm and a height of 6 cm. Each solid block will have a volume of 360 cm3. How many cylinders of gold will it take to mold one block of gold?

OpenStudy (campbell_st):

start by finding the radius of the cylinder. then volume of a cylinder is \[V = \pi \times r^2\times h\] r = radius and h = height substitute and find the volume of gold in the cylinder. the last part is the number of cylinders. it will be volume of the block divided by volume of a cylinder. hope this helps

OpenStudy (anonymous):

Thanks this helps alot